# CERN-TH-2023-223 HERWIG-2023-01 KA-TP-28-2023

### v1

@@ 1 Introduction
Herwig, a multi-purpose event generator for reactions at $pp$, $ep$ and $ee$ colliders, is now available in a new version, $7.3$. The Herwig 7 release series, started with Bellm:2015jjp , is based on the Herwig++ development Bahr:2008pv ; Bahr:2008tx ; Bahr:2008tf ; Gieseke:2011na ; Arnold:2012fq ; Bellm:2013lba and fully supersedes the previous Herwig++ and HERWIG versions. The main cornerstones of the Herwig 7 series are significant improvements to the prediction of the hard scattering, now routinely available at next-to-leading order QCD, as well as major development and theoretical insight into the available angular-ordered and dipole parton shower algorithms and their accuracy. A major follow-up release of these previous Herwig 7 versions is now available as Herwig 7.3. The updates in this release are centred around parton shower algorithms, the hadronization model and soft QCD interactions in hadronic collisions. The current release note also describes features which will subsequently appear as minor digit Herwig 7.3.x releases in the near future. Please refer to the Herwig++ manual Bahr:2008pv , the Herwig 7.0 Bellm:2015jjp as well as this release note when using the new version of the program. Studies or analyses that rely on a particular feature of the program should also reference the paper(s) where the physics of that feature was first described. The authors are happy to provide guidance on which features are relevant for a particular analysis. A major publication intended as physics reference and manual will be appearing in the coming year.
@@ 3.1 HQET in hadronization and decay of heavy and excited hadrons
could result in heavy quarks acting as static color sources. This facilitates the emergence of a “spin-flavour symmetry” for heavy quarks. Meanwhile, the Falk-Peskin “no-win” theorem Falk:1993rf elucidates the limits of polarization information retrieval:
@@
At this stage, it is crucial to note that due to the scarce experimental data around these heavy hadrons, we depend heavily on the principles of HQET to predict how they decay, including their lifetimes and decay modes, as it lays down fundamental connections between the coupling constants involved in these decays. HQET has been already used for modeling the strong and radiative breakdown of heavy baryons, as shown in some prior studies and its integration into Herwig++ Bahr:2008pv . Nevertheless, we still have to develop such a prescription for the heavy meson sector. We begin by identifying the $s-$ and $p-$wave meson multiplets: the ground state doublets ($J^{P}=0^{-},1^{-}$) including $D$ and $D^{\star}$ mesons, the $J^{P}=1^{+},2^{+}$ doublet with $D_{1}$ and $D_{2}^{\star}$ mesons, and the $J^{P}=0^{+},1^{+}$ doublet comprising $D_{0}^{\star}$ and $D_{1}^{\prime}$ mesons Falk:1992cx . We also account for the mixing effects between $D_{1}$ and $D_{1}^{\prime}$ mesons resulting from higher-order corrections in HQET. The methodology and notation for calculating the matrix elements associated with meson decays are as per Falk et al. Falk:1992cx ; Falk:1995th , with a focus on balancing theoretical precision with empirical data representation. The matrix elements are defined by equations (20) through (25) below, incorporating decay parameters and polarization vectors, where $p_{i}$ denotes the hadron momenta and $\epsilon_{i}$ the polarization vectors LE_PAPER . Here, the index $i=0$ identifies the parent hadron, while $i=1$ and $i=2$ refer to the resulting heavy and light daughter hadrons. The mass of a hadron $H$ is represented by $m_{H}$. The parameters $g$, $h$, $\Lambda$, $f_{\pi}$ and $f^{\prime\prime}$ are constants associated with the decay process. Furthermore, $\theta_{q}$ is introduced as the mixing angle between the meson pairs ($D_{1}$, $D^{\prime}_{1}$) and ($D_{s1}$, $D^{\prime}_{s1}$). {strip}
@@
To integrate the processes of strong and radiative decays for excited heavy mesons into Herwig 7, we have created two dedicated classes: HQETStrongDecayer and HQETRadiativeDecayer. The decay parameters, emphasized in this segment, are set as variables that users can modify. Such an arrangement offers adaptability, allowing for adjustments and enhancements in line with new findings or needs. The efficacy of using HQET and spin-flavor symmetry to predict polarization-sensitive measurements has been extensively discussed in LE_PAPER , see e.g. Figure 3.
@@ 3.3 New infrastructure for hadronization models
Several improvements are in development with respect to the cluster hadronization and will soon be available with an upcoming Herwig release, among them a dynamic cluster model DynamicHadronizationPaper and other options for an improved understanding of hadronization. Many of the structures already appear in this release and allow for the flexible adjustment of kinematics and mass distributions in the cluster fission process (offering more alternatives to the choices made in Sec. 3.2), as well as the generation of dynamic gluon masses using the GluonMassGenerator class.

### v2

@@ 1 Introduction
Herwig, a multi-purpose event generator for reactions at $pp$, $ep$ and $ee$ colliders, is now available in a new version, $7.3$. The Herwig 7 release series, started with Bellm:2015jjp and continuing with Bellm:2017bvx ; Bellm:2019zci , is based on the Herwig++ development Bahr:2008pv ; Bahr:2008tx ; Bahr:2008tf ; Gieseke:2011na ; Arnold:2012fq ; Bellm:2013lba and fully supersedes the previous Herwig++ and HERWIG versions. The main cornerstones of the Herwig 7 series are significant improvements to the prediction of the hard scattering, now routinely available at next-to-leading order QCD, as well as major development and theoretical insight into the available angular-ordered and dipole parton shower algorithms and their accuracy. A major follow-up release of these previous Herwig 7 versions is now available as Herwig 7.3. The updates in this release are centred around parton shower algorithms, the hadronization model and soft QCD interactions in hadronic collisions. Please refer to the Herwig++ manual Bahr:2008pv , the Herwig 7.0 Bellm:2015jjp as well as this release note when using the new version of the program. Studies or analyses that rely on a particular feature of the program should also reference the paper(s) where the physics of that feature was first described. The authors are happy to provide guidance on which features are relevant for a particular analysis. A major publication intended as physics reference and manual will be appearing in the coming year.
@@ 3.1 HQET in hadronization and decay of heavy and excited hadrons
could result in heavy quarks acting as static colour sources. This facilitates the emergence of a “spin-flavour symmetry” for heavy quarks. Meanwhile, the Falk-Peskin “no-win” theorem Falk:1993rf elucidates the limits of polarization information retrieval:
@@
At this stage, it is crucial to note that due to the scarce experimental data around these heavy hadrons, we depend heavily on the principles of HQET to predict how they decay, including their lifetimes and decay modes, as it lays down fundamental connections between the coupling constants involved in these decays. HQET has been already used for modeling the strong and radiative decay of heavy baryons, as shown in some prior studies and its integration into Herwig++ Bahr:2008pv . Nevertheless, we still have to develop such a prescription for the heavy meson sector. We begin by identifying the $s-$ and $p-$wave meson multiplets: the ground state doublets ($J^{P}=0^{-},1^{-}$) including $D$ and $D^{\star}$ mesons, the $J^{P}=1^{+},2^{+}$ doublet with $D_{1}$ and $D_{2}^{\star}$ mesons, and the $J^{P}=0^{+},1^{+}$ doublet comprising $D_{0}^{\star}$ and $D_{1}^{\prime}$ mesons Falk:1992cx . We also account for the mixing effects between $D_{1}$ and $D_{1}^{\prime}$ mesons resulting from higher-order corrections in HQET. The methodology and notation for calculating the matrix elements associated with meson decays are as per Falk et al. Falk:1992cx ; Falk:1995th , with a focus on balancing theoretical precision with empirical data representation. The matrix elements are defined by equations (20) through (25) below, incorporating decay parameters and polarization vectors, where $p_{i}$ denotes the hadron momenta and $\epsilon_{i}$ the polarization vectors LE_PAPER . Here, the index $i=0$ identifies the parent hadron, while $i=1$ and $i=2$ refer to the resulting heavy and light daughter hadrons. The mass of a hadron $H$ is represented by $m_{H}$. The parameters $g$, $h$, $\Lambda$, $f_{\pi}$ and $f^{\prime\prime}$ are constants associated with the decay process. Furthermore, $\theta_{q}$ is introduced as the mixing angle between the meson pairs ($D_{1}$, $D^{\prime}_{1}$) and ($D_{s1}$, $D^{\prime}_{s1}$). {strip}
@@
To integrate the processes of strong and radiative decays for excited heavy mesons into Herwig 7, we have created two dedicated classes: HQETStrongDecayer and HQETRadiativeDecayer. The decay parameters, emphasized in this segment, are set as variables that users can modify. Such an arrangement offers adaptability, allowing for adjustments and enhancements in line with new findings or needs. The efficacy of using HQET and spin-flavour symmetry to predict polarization-sensitive measurements has been extensively discussed in LE_PAPER , see e.g. Figure 3.
@@ 3.3 New infrastructure for hadronization models
Several improvements are in development with respect to the cluster hadronization and will soon be available with an upcoming Herwig release, among them a dynamic cluster model Hoang:2024zwl and other options for an improved understanding of hadronization. Many of the structures already appear in this release and allow for the flexible adjustment of kinematics and mass distributions in the cluster fission process (offering more alternatives to the choices made in Sec. 3.2), as well as the generation of dynamic gluon masses using the GluonMassGenerator class.